# Old Meets New: Connecting Two Infinite Families of Congruences Modulo Powers of 5 for Generalized Frobenius Partition Functions

### v1

@@ Abstract.
In 2012 Paule and Radu proved a difficult family of congruences modulo powers of 5 for Andrews’ 2-colored generalized Frobenius partition function. The family is associated with the classical modular curve of level 20. We demonstrate the existence of a second congruence family for a related generalized Frobenius partition function associated with the same curve. Unlike traditional proofs, we construct an isomorphism between this new family and the original family of congruences via an automorphism on the associated ring of functions. We also give some important insights into the behavior of these congruence families with respect to the Atkin–Lehner involution which proved very important in Paule and Radu’s original proof.
@@ 1.1. Background
In 1994 the first author proposed the following congruence family for the generalized 2-colored Frobenius partition function $c\phi_{2}(n)$:
@@
What makes this result so interesting is that, while it superficially resembles the classical congruence families of the unrestricted partition function $p(n)$ modulo powers of 5, 7, 11 that Ramanujan first discovered [10], it is a much more difficult result to prove. Indeed, a proof was not found until Paule and Radu’s work in 2012 [7], 18 years after its proposal, and some 93 years after Ramanujan founded the subject.
@@
The second author has invented some more general techniques, embodied in the localization method, in an attempt to give a unified framework to the subject of partition congruences. These techniques work in many different cases (e.g., [11], [12]), and indeed reduce to standard methods for classical congruence families. However, the congruences of Theorem 1.1 are strangely resistant to localization. As such, we cannot at present extend classical methods via a systematic approach to congruence families of a certain class which encompass Theorem 1.1 and other contemporary congruence families, some of which are yet standing conjectures. In a sense, we know that the family is correct, but we still do not fully understand why it is there, and why more systematic approaches fail to resolve it.
In this work we report on recent progress in studying the underlying causes of this family. To our surprise, we have discovered that Theorem 1.1 is but part of a natural and extremely closely related pair of congruence families, exhibited by a pair of Frobenius partition functions. It may be that one reason that the congruences for $c\phi_{2}$ have proved so difficult to understand is that in the decade since they were resolved (and the three decades since they were first proposed), they have been effectively half perceived.
@@
The latter theorem already suggests a close relationship between the two families. However, we were surprised to discover just how close this connection is. In the first place, each family is associated with a sequence of functions defined on the classical modular curve $\mathrm{X}_{0}(20)$—a curve of genus 1 and cusp count 6. For each case, one constructs a pair of rank 2 $\mathbb{Z}[t]$-modules (with $t$ a natural Hauptmodul corresponding to $\mathrm{X}_{0}(5)$), and demonstrates membership of the respective functions within these modules.
For each family, the generators of the respective modules are different; however, we have found that the proof of Theorem 1.2 is—at least in terms of formal algebraic manipulations—identical to that of Theorem 1.1. Indeed, both congruence families are related to each other via a certain automorphism between subrings of the modular functions on $\mathrm{X}_{0}(20)$ which fixes the modular functions on $\mathrm{X}_{0}(5)$. This may have enormous implications on the future study of congruence families associated with modular curves of composite level.
@@
Finally, these families bear a strange symmetry to one another upon the application of a certain Atkin–Lehner involution to their respective functions. This symmetry gives rise to an absolutely astonishing and beautiful result which explains a curious coefficient behavior that was discovered by Paule and Radu in their original work [7, Section 6], and used to assist in the construction of their proof of Theorem 1.1.
The remainder of our paper is organized as follows: in Section 2 we give the definitions for the associated Frobenius partitions as well as the generating functions relevant to our work. In Section 3 we give the underlying modular function spaces over $\mathrm{X}_{0}(20)$ that will be relevant to us. We prove the initial stages of both congruence families via the associated initial terms on the related function sequences. We then show how Theorem 1.2 can be derived from Theorem 1.1 via a certain automorphism on a subring of modular functions of $\mathrm{X}_{0}(20)$.
In Section 4 we show how these two sequences, when combined, produce functions which reduce to the simpler curve $\mathrm{X}_{0}(10)$. We then give conjectures as to the accessibility of proving the two congruence families relative to one another via more systematic techniques.
@@
In this work we presume an understanding of the theory of modular functions and its associated modular curves. See [11, Section 2]. A useful classical reference is [6], and a more modern reference is [4, Chapters 2-3].
@@ 2. Frobenius Partition Functions
  • Each $a_{i},b_{j}\in\mathbb{Z}_{\geq 0}$, and any integer occurs at most $k$ times,
@@
is the standard $U_{5}$ operator (see [6, Chapter 8] for a list of the standard properties of $U_{5}$), and
@@
It is a comparatively standard procedure in the theory to show that for all $\alpha\geq 1$ we have
@@
This is the Hauptmodul for $\mathrm{X}_{0}(5)$, a curve of genus 0 and cusp count 2. Indeed, $t$ is the critical function for proving Ramanujan’s congruence family for $p(n)$ modulo powers of 5 (see [14] or [6, Chapters 7-8]). Being modular over $\Gamma_{0}(5)$, it is clearly modular over $\Gamma_{0}(20)$. It does not live at the cusp $[0]$, but we can perform the same trick we did with $L_{1}^{(1)}$—we can represent $t$ as a rational polynomial. Indeed, we have
@@
Thus through an elementary set of manipulations, we can now represent elements of our localization ring as polynomials in $t$, $x$, $y$. Indeed, one sees that $x^{3}$ can be written as a polynomial in $t$, $x$, $x^{2}$. As such, we can express each $L_{\alpha}^{(\beta)}$ as a polynomial in $t,x,y$, but with the powers of $x$ restricted to 2 or lower (and the powers of $y$ restricted to 1 or lower). Thus, our representations which will be “mostly” in terms of the function $t$.
@@
Comparing (3.22) and (3.28), one finds that the two identifies formally match. Indeed, we have verified that the proofs, as well as the representations themselves, formally match exactly—they differ only in the exact form of the module generators $p^{(\beta)}_{\alpha}$.
@@
$\mathcal{R}:=\mathbb{C}[t]+p^{(0)}_{0}\mathbb{C}[t]\ +$ $p^{(0)}_{1}\mathbb{C}[t]+p^{(1)}_{0}\mathbb{C}[t]+p^{(1)}_{1}\mathbb{C}[t],$ (3.29)
and the ring automorphisms
$\sigma^{(\beta)}$ $:\mathcal{R}\longrightarrow\mathcal{R}$ (3.30)
@@
This theorem is proved by examining the action of $U^{(\beta)}_{a}$ on $t^{m}$ and $t^{m}p^{(\beta)}_{a}$. That is, one simply must show that $U^{(0)}_{a}\left(t^{m}\right)$ (resp. $U^{(0)}_{a}\left(t^{m}p^{(0)}_{a}\right)$) formally matches $U^{(1)}_{a}\left(t^{m}\right)$ (resp. $U^{(1)}_{a}\left(t^{m}p^{(1)}_{a}\right)$) for $a=0,1$; the only difference is switching out one generator ($p^{(0)}_{a}$) for another ($p^{(1)}_{a}$).
In particular, $t$ is subject to a modular equation of degree 5 which is not dependent on our module generators [7, Definition 3.1, Lemma 3.2]. With this equation we can build a recurrence of $U^{(\beta)}_{a}\left(t^{m}\right)$, $U^{(\beta)}_{a}\left(t^{m}p^{(0)}_{a}\right)$ for any $m$, provided that we have precise expressions for 5 consecutive values of $m$ [7, Lemma 3.3].
In our Mathematica supplement (see the end of Section 1) we derive the explicit relations of $U^{(0)}_{a}\left(t^{m}p^{(0)}_{a}\right)$ for $a=0,1$, and five consecutive values of $m$. We then show that they match the relations of $U^{(1)}_{a}\left(t^{m}p^{(0)}_{a}\right)$ for the same values of $a$, $m$, which are derived by Paule and Radu in [7, Section 7].
In our case, we use the same values of $a$, $m$ that Paule and Radu used to expedite the comparison. We can prove these relations with the modular cusp analysis. This is also done in our Mathematica supplement (see the end of Section 1).

### v2

@@ Abstract.
In 2012 Paule and Radu proved a difficult family of congruences modulo powers of 5 for Andrews’ 2-colored generalized Frobenius partition function. The family is associated with the classical modular curve of level 20. We demonstrate the existence of a congruence family for a related generalized Frobenius partition function associated with the same curve. Unlike traditional proofs, we construct an isomorphism between this new family and the original family of congruences via a mapping on the associated rings of modular functions. The existence of this second congruence family and its relationship with the first poses significant implications for future work on congruences associated with modular curves of composite level. We show how a similar approach can be made to multiple other recent examples in the literature. We also give some important insights into the behavior of these congruence families with respect to the Atkin–Lehner involution which proved very important in Paule and Radu’s original proof.
@@ 1. Introduction
In 1994 the second author proposed the following congruence family for the generalized 2-colored Frobenius partition function $c\phi_{2}(n)$:
@@
This congruence family deceptively resembles the classic congruences for the partition function $p(n)$ discovered by Ramanujan, but presents much greater challenges; a complete proof was not found until 2012.
Our principal discovery is that the congruence family of Theorem 1.1 is only one specific manifestation of a more general $\ell$-adic behavior on sequences of modular functions on the curve $\mathrm{X}_{0}(20)$. In particular, this behavior manifests itself as a second congruence family for the Fourier coefficients of a different modular form.
It appears that a remarkable property exists for congruence families associated with modular curves of composite level—in some sense, the same family can manifest in the Fourier coefficients associated with multiple different modular forms. We show how our approach may be applied to many other recently studied partition functions.
We also discuss some of the implications of this approach to the localization method of congruence families, and some peculiar properties that the functions associated with Theorem 1.1 have with respect to Ramanujan’s phi function.
1.1. Background
What makes Theorem 1.1 so interesting is that, while it superficially resembles the classical congruence families of the unrestricted partition function $p(n)$ modulo powers of 5, 7, 11 that Ramanujan first discovered [18], it is a much more difficult result to prove. Indeed, a proof was not found until Paule and Radu’s work in 2012 [15], 18 years after its proposal, and some 93 years after Ramanujan founded the subject.
@@ 1.1. Background
The third author has invented some more general techniques, embodied in the localization method, in an attempt to give a unified framework to the subject of partition congruences. These techniques work in many different cases (e.g., [19], [20]), and indeed reduce to standard methods for classical congruence families. However, the congruences of Theorem 1.1 are strangely resistant to localization. As such, we cannot at present extend classical methods via a systematic approach to congruence families of a certain class which encompass Theorem 1.1 and other contemporary congruence families, some of which are yet standing conjectures. In a sense, we know that the family is correct, but we still do not fully understand why it is there, and why more systematic approaches fail to resolve it.
In this work we report on recent progress in studying the underlying difficulties of this family. To our surprise, we have discovered that Theorem 1.1 is but part of a natural and extremely closely related pair of congruence families, exhibited by a pair of Frobenius partition functions. It may be that one reason that the congruences for $c\phi_{2}$ have proved so difficult to understand is that in the decade since they were resolved (and the three decades since they were first proposed), they have been effectively half perceived.
@@
The latter corollary already suggests a close relationship between the two families. However, we were surprised to discover just how close this connection is. In the first place, each family is associated with a sequence of functions defined on the classical modular curve $\mathrm{X}_{0}(20)$—a curve of genus 1 and cusp count 6. For each case, one constructs a pair of rank 2 $\mathbb{Z}[t]$-modules (with $t$ a natural Hauptmodul corresponding to $\mathrm{X}_{0}(5)$), and demonstrates membership of the respective functions within these modules.
For each family, the generators of the respective modules are different; however, we have found that the proof of Theorem 1.2 is—at least in terms of formal algebraic manipulations—identical to that of Theorem 1.1. Indeed, the families are equivalent; they are related to each other via a certain invertible mapping between subrings of the modular functions on $\mathrm{X}_{0}(20)$ which fixes the modular functions on $\mathrm{X}_{0}(5)$.
It appears that in some strange sense Theorems 1.1 and 1.2 constitute a “single” congruence family (i.e., a specific pattern of $5$-adic convergence to 0), which somehow manifests itself in two different ways, for two different functions. This approach may have enormous implications on the future study of congruence families associated with modular curves of composite level.
@@
Finally, these families bear a strange symmetry to one another upon the application of a certain Atkin–Lehner involution to their respective functions. This symmetry gives rise to an absolutely astonishing and beautiful result which explains a curious coefficient behavior that was discovered by Paule and Radu in their original work [15, Section 6], and used to assist in the construction of their proof of Theorem 1.1.
The remainder of our paper is organized as follows: in Section 2 we give the definitions for the associated Frobenius partitions as well as the generating functions relevant to our work. In Section 3 we give the underlying modular function spaces over $\mathrm{X}_{0}(20)$ that will be relevant to us. We prove the initial stages of both congruence families via the associated initial terms on the related function sequences. We then show how Theorem 1.2 can be derived from Theorem 1.1 via a certain mapping on a subring of modular functions of $\mathrm{X}_{0}(20)$. We also give an explicit form for this mapping, namely as the action of a specific element of $\Gamma_{0}(10)$.
In Section 4 we describe more examples of partition congruences in which a similar approach can be used to show that certain divisibility properties are shared between functions which at first sight appear very different. The first is a pair of congruences exhibited by the Frobenius partition functions $\phi_{2}(n)$ and $\psi_{2}(n)$, studied by Andrews [1] and by Eckland and the second author [10]. Individual congruences are rarer for these functions, and no infinite families are known to exist. Our method suggests that any families for one of these functions would necessarily imply an associated family for the other. Moreover, we show how a witness identity for a congruence for $\phi_{2}(n)$, which is derived by a comparatively tedious cusp analysis, can be used to derive (in a far less tedious manner) an analogous witness identity for the related congruence for $\psi_{2}(n)$.
The second example involves a pair of altogether different functions sudied by Connor Morrow in collaboration with the first author [11]. One is a congruence family exhibited by a function related to the crank parity function originally proved by Choi, Kang, and Lovejoy in [7]. The other is a family exhibited by a function closely related to the 1-shell TSPP and $\overline{\mathcal{EO}}(n)$ functions studied in [2], [5], and [12]. Moreover, we note that yet more extremely recent work in this subject has been done by the first author in concert with Dandan Chen and Rong Chen [4].
Still another striking example involves a remarkable pair of infinite internal congruence families exhibited by coefficients related to quotients of Ramanujan’s theta functions, studied by Chern and Tang [6].
Note the variety of different functions in our examples. We have Frobenius partitions, of course; but we also have objects related to mock theta functions, the partition crank statistic, and Ramanujan’s classic theta functions. Other potential examples that warrant investigation include smallest parts functions (especially those related to mock theta functions), $k$-elongated plane partitions, and the Hecke trace functions $t_{m}(d)$.
The key unifying property of all of these examples is that the modular curve associated with the congruence families has composite level. The latter property also appears to be critical in studying the most difficult problems of the subject.
As such, a proper understanding of “equivalent congruences” associated with composite level modular curves appears to be an extremely important phenomenon, though its universality has not hitherto been fully grasped. We hope that we have effectively conveyed the significance of this approach to future work.
In Section 5 we show how the two function sequences corresponding to Theorems 1.1 and 1.2, when combined, produce functions which reduce to the simpler curve $\mathrm{X}_{0}(10)$. This has implications for the more systematic programmes (e.g. localization) which have hitherto failed to account for congruence families like those of Theorems 1.1 and 1.2. We give conjectures as to the accessibility of proving the two congruence families relative to one another via localization.
@@
In this work we presume an understanding of the theory of modular functions and its associated modular curves. See [19, Section 2]. A useful classical reference is [14], and a more modern reference is [8, Chapters 2-3].
@@ 2. Frobenius Partition Functions
  • Each $a_{i},b_{j}$ belongs to one of $k$ copies of $\mathbb{Z}_{\geq 0}$,
@@
Here we use the standard $q$-Pochhammer symbol
$(q^{a};q^{b})_{\infty}:=\prod_{m=0}^{\infty}\left(1-q^{bm+a}\right).$
@@
is the standard $U_{5}$ operator (see [14, Chapter 8] for a list of the standard properties of $U_{5}$), and
@@
It is a comparatively standard procedure in the theory to show that for all $\alpha\geq 1$, with $a\in\{0,1\}$ matching the parity of $\alpha$, we have
@@
This is the Hauptmodul for $\mathrm{X}_{0}(5)$, a curve of genus 0 and cusp count 2. Indeed, $t$ is the critical function for proving Ramanujan’s congruence family for $p(n)$ modulo powers of 5 (see [22] or [14, Chapters 7-8]). Being modular over $\Gamma_{0}(5)$, it is clearly modular over $\Gamma_{0}(20)$. It does not live at the cusp $[0]$, but we can perform the same trick we did with $L_{1}^{(1)}$—we can represent $t$ as a rational polynomial. Indeed, we have
@@
Thus through an elementary set of manipulations, we can now represent elements of our localization ring as polynomials in $t$, $x$, $y$. Indeed, one sees that $x^{3}$ can be written as a polynomial in $t$, $x$, $x^{2}$. As such, we can express each $L_{\alpha}^{(\beta)}$ as a polynomial in $t,x,y$, but with the powers of $x$ restricted to 2 or lower (and the powers of $y$ restricted to 1 or lower). Thus, our representations will be “mostly” in terms of the function $t$.
@@
Comparing (3.22) and (3.28), one finds that the two identities formally match. Indeed, we have verified that all of the formal representations for $L_{\alpha}^{(0)}$ and $L_{\alpha}^{(1)}$ match exactly—they differ only in the associated module generators $p^{(\beta)}_{\alpha}$.
@@
$\mathcal{R}^{(1)}:$ $=\mathbb{C}(t)+p^{(1)}_{0}\mathbb{C}(t)\ +p^{(1)}_{1}\mathbb{C}(t)$ (3.29)
$\mathcal{R}^{(0)}:$ $=\mathbb{C}(t)+p^{(0)}_{0}\mathbb{C}(t)\ +p^{(0)}_{1}\mathbb{C}(t)$ (3.30)
and the ring mappings
$\sigma^{(\beta)}$ $:\mathcal{R}^{(\beta)}\longrightarrow\mathcal{R}^{(1-\beta)}$ (3.31)
@@
3.1. Proof of Theorem 3.1
One straightforward, if tedious, proof of Theorem 3.1 involves directly computing and comparing $U^{(\beta)}_{a}\left(t^{m}\right)$, $U^{(\beta)}_{a}\left(t^{m}p^{(0)}_{a}\right)$ for $\beta=0,1$, $a=0,1$, and $m$ ranging over 5 consecutive values. One can then invoke a modular equation for $t$, e.g., the equation used by Paule and Radu in [15].
However, the existence of such a ring mapping system as $\sigma^{(\beta)}$ strongly suggests that there exists some explicit means of transforming the functions corresponding to $c\psi_{2,0}(n)$ into those corresponding to $c\psi_{2,1}(n)$. Our first instinct on the matter was to attempt a connection between $\mathrm{C}\Psi_{2,\beta}$, $\mathrm{C}\Psi_{2,1-\beta}$ via an Atkin–Lehner involution.
At first no direct connection could be found. On closer examination, the first author realized that we can indeed create an explicit mapping between these generating functions by first mapping $q\mapsto-q$, then applying an Atkin–Lehner involution, and finally mapping $q\mapsto-q$ a second time. Equivalently, we conjugate a certain Atkin–Lehner involution by a half-integer valued matrix which equates to sending $q\mapsto-q$.
More specifically, we consider the operator
$W:=\begin{pmatrix}4&-1\\ 100&-24\end{pmatrix}.$
Let us start by taking the functions $\mathcal{A}^{(\beta)}$ which are necessary for construction of our $U^{(\beta)}_{\alpha}$ operators. If we take, say, $\beta=1$, and we first map $q\mapsto-q$, then we have
$\mathcal{A}^{(1)}\longmapsto\frac{\eta(\tau)^{4}\eta(4\tau)^{2}\eta(50\tau)^{7}}{\eta(2\tau)^{7}\eta(25\tau)^{4}\eta(100\tau)^{2}}.$ (3.34)
We want to send $\tau$ to $W\tau$. See Section 6 below for a detailed treatment of how $W$ affects each eta factor, and indeed how it affects our relevant spaces of modular functions. This gives us
$\frac{\eta(\tau)^{4}\eta(4\tau)^{2}\eta(50\tau)^{7}}{\eta(2\tau)^{7}\eta(25\tau)^{4}\eta(100\tau)^{2}}\longmapsto\frac{\eta(\tau)^{2}\eta(4\tau)^{4}\eta(50\tau)^{7}}{\eta(2\tau)^{7}\eta(25\tau)^{2}\eta(100\tau)^{4}}.$ (3.35)
Finally, applying $q\mapsto-q$, we achieve
$\frac{\eta(\tau)^{2}\eta(4\tau)^{4}\eta(50\tau)^{7}}{\eta(2\tau)^{7}\eta(25\tau)^{2}\eta(100\tau)^{4}}\longmapsto\mathcal{A}^{(0)}.$ (3.36)
We now have a direct means of connecting the two.
If we examine the composition of these mappings, we find that the matrix representing the overall action is
$\begin{pmatrix}1&1/2\\ 0&1\end{pmatrix}W\begin{pmatrix}1&1/2\\ 0&1\end{pmatrix}=\begin{pmatrix}54&14\\ 100&26\end{pmatrix},$ (3.37)
which of course will reduce via the associated transformation to
$\gamma=\begin{pmatrix}27&7\\ 50&13\end{pmatrix}\in\Gamma_{0}(10).$
Notice that by this transformation, we immediately have
$t\left(\gamma\tau\right)=t(\tau),$
$x\left(\gamma\tau\right)=x(\tau),$
as the former is modular over $\Gamma_{0}(5)$, and the latter is modular over $\Gamma_{0}(10)$. We need only to understand how $\gamma$ affects $y$. Applying the transformation and accounting for the Nebentypus factor, we can show that
$y\left(\gamma\tau\right)=-\frac{1}{4}(1+6x+4y).$
Knowing that $t$ and $x$ are invariant, we can make this substitution for $y$ into $p_{1}^{(1)}$, one gets
$p^{(1)}_{1}\left(\gamma\tau\right)=$ $\ t+25t^{2}-8tx-y\left(\gamma\tau\right)\left(4+8t+500t^{2}+32x-480tx+64x^{2}-1600tx^{2}\right)$
$=$ $\ 1+9t+8x+64tx+16x^{2}+80tx^{2}+4y+8ty+500t^{2}y+32xy-480txy+64x^{2}y-1600tx^{2}y$
$=$ $\ p^{(0)}_{1}.$
Similarly, we have
$p^{(1)}_{0}\left(\gamma\tau\right)=$ $\ p^{(0)}_{0}.$
Thus $\sigma^{(1)}$ equates to sending $f(\tau)$ to $f(\gamma\tau)$, and we can define
$\sigma^{(1)}\left(f\right)=f\left(\gamma\tau\right).$
We of course have
$L_{1}^{(0)}$ $=U_{5}\left(\mathcal{A}^{(0)}\right)=U_{5}\left(\mathcal{A}^{(1)}(\gamma\tau)\right)=L_{1}^{(1)}(\gamma\tau)=\sigma^{(1)}\left(L_{1}^{(1)}\right).$ (3.38)
Supposing that for some $\alpha\geq 2$ we have $L_{\alpha-1}^{(0)}=L_{\alpha-1}^{(1)}(\gamma\tau)$. Again allowing $\alpha\equiv a\pmod{2}$ for $a\in\{0,1\}$, we have
$L_{\alpha}^{(0)}$ $=U_{1-a}^{(0)}\left(L_{\alpha-1}^{(0)}\right)$ (3.39)
$=U_{5}\left(\left(\mathcal{A}^{(0)}\right)^{a}L_{\alpha-1}^{(0)}\right)$ (3.40)
$=U_{5}\left(\left(\mathcal{A}^{(1)}(\gamma\tau)\right)^{a}L_{\alpha-1}^{(1)}(\gamma\tau)\right)$ (3.41)
$=U_{5}\left(\left(\mathcal{A}^{(1)}\right)^{a}L_{\alpha-1}^{(1)}\right)(\gamma\tau)$ (3.42)
$=L_{\alpha}^{(1)}(\gamma\tau)$ (3.43)
$=\sigma^{(1)}\left(L_{\alpha}^{(1)}\right).$ (3.44)
Ordinarily, we could not justify (3.41)-(3.42). However, in our case the action of $\gamma$ is equivalent to the composition of transformations in (3.37), and the factors $W$, $\left(\begin{smallmatrix}1&1/2\\ 0&1\end{smallmatrix}\right)$ commute with the $U_{5}$ operator.
Thus induction on $\alpha$ gives us
$\sigma^{(1)}\left(L_{\alpha}^{(1)}\right)=L_{\alpha}^{(1)}\left(\gamma\tau\right)=L_{\alpha}^{(0)}$
for all $\alpha\geq 1$, whereupon we have Theorem 3.1.